Leonidas the Lion started at one end of a meadow and strolled 3/4 of its length. Zeus the Zebra started at the opposite end of the same meadow and crawled 2/3 of its length. What fraction of the length of the meadow are Leonidas and Zeus now apart?
Answer by greenestamps(13200)



Picture the two ends of the meadow as 0 and 1 on a number line.

Leonidas starts at 0 and moves 3/4 of the way towards 1, ending up at 0+3/4 = 3/4.

Zeus starts at 1 and moves 2/3 of the way towards 0, ending up at 1-2/3 = 1/3.

The fraction of the meadow between them is the fraction between 3/4 and 1/3.

3/4 - 1/3 = ______